OpenStudy (anonymous):

Help me with this please. Find a point on the curve y=x^2 that is closest to the point (0,2).

OpenStudy (anonymous):

ok just the distance formula, or rather its square

OpenStudy (anonymous):

you have a point (0,2) and a point on the curve (x,x^2) and the distance between two is \[D(x)=(x-0)^2+(x^2-2)^2\]

OpenStudy (anonymous):

ok that is not really the distance, because we didn't use the square root, so it is the square of the distance. you do get to use calculus right?

OpenStudy (nilankshi):

help me also

OpenStudy (anonymous):

\[D(x)=x^4-3x^2+2\] if my algebra is correct, (it is early) and you want the minimum. take the derivative, set = 0 and solve \[D'(x)=4x^3-6x\] put \[4x^3-6x=0\] \[2x(2x-3)=0\] so choices are \[x=0\] or \[x=\frac{3}{2}\]

OpenStudy (anonymous):

clearly 0 works because it is the same point. so i guess you are supposed to say \[(\frac{3}{2},\frac{9}{4})\] again it is early so make sure to check my work

OpenStudy (amistre64):

i look at this way\[\]the shortest distance is a line from the point that make a perpendicular to the tangent of the curve

OpenStudy (anonymous):

good morning. how did i do?

OpenStudy (amistre64):

mornin, i see you survived the aftershocks aswell :)

OpenStudy (anonymous):

yes and the tsunami

OpenStudy (amistre64):

y=x^2; y' = 2x defines our slope; now we need to develop a notion for the slope from the point (0,2) to a given point on the line y = x^2, or rather: (x,x^2)\[\]

OpenStudy (amistre64):

are the lines working

OpenStudy (nilankshi):

see my qus also

OpenStudy (nilankshi):

pls help me

OpenStudy (amistre64):

(x,x^2) -(0, 2) -------- (x, x^2-2); slope =\(\cfrac{x^2-2}{x}\), right?

OpenStudy (amistre64):

grrr.....

OpenStudy (anonymous):

uhm...I'll process this first, but yeah we're using calculus.:))..Thanks for the help SATELLITE 73..I got disconnected. And thank you amistre64. I understand now. :)

OpenStudy (amistre64):

per slope is flip and negative right?

OpenStudy (anonymous):

hmm i am following but getting a different equation than the one i solved.

OpenStudy (amistre64):

2x = \(-\cfrac{x}{x^2-2}\), and solve for x is my thought

OpenStudy (anonymous):

oh i see my dumb mistake when i was using your method. yea this will do it and get the exact same equation too right?

OpenStudy (amistre64):

2 = -1/(x^2-2)

OpenStudy (amistre64):

x^2-2=-1/2

OpenStudy (anonymous):

\[2x^3-4x=-x\] \[2x^3-3x=0\] etc

OpenStudy (amistre64):

x^2 = 1.5

OpenStudy (amistre64):

x= sqrt(1.5); maybe?

OpenStudy (anonymous):

ok now we diverge

OpenStudy (amistre64):

its prolly my brain math

OpenStudy (anonymous):

\[2x=-\frac{x}{x^2-2}\] \[2x^3-4x=-x\] \[2x^2-3x=0\] \[x(2x-3)=0\] \[x=\frac{3}{2}\]

OpenStudy (amistre64):

i divided off the xs first

OpenStudy (anonymous):

hmmm well you lose a root that way but lets see what happens, because there is more than one way to skin a moose

OpenStudy (amistre64):

x = sqrt(3/2)

OpenStudy (anonymous):

now i am thoroughly confused!

OpenStudy (amistre64):

\[2x = \frac{-x}{x^2-2}\] \[2 = \frac{-1}{x^2-2}\] \[2(x^2-2) = -1\] \[2x^2-4+1 = 0\] \[2x^2-3 = 0\] \[x^2 = \frac{3}{2}\]

OpenStudy (anonymous):

i cleared the fraction and got \[2x^3-4x=-x\] or \[2x^3-3x=0\] which made me happy because it is the same equation i got form the distance formula

OpenStudy (anonymous):

oh wait i see

OpenStudy (anonymous):

you cannot divide by x!

OpenStudy (amistre64):



OpenStudy (amistre64):

becasue x could be zero roght?

OpenStudy (anonymous):

my professor's answer was y=3/2 , x=+- (sq rt of 3/2)does this make sense?

OpenStudy (anonymous):

yes it makes sense and i am a moron

OpenStudy (anonymous):

i was solving for y and got \[\frac{3}{2}\]

OpenStudy (amistre64):

when you find the points of interest; we plug them in to see which ones work

OpenStudy (amistre64):

-sqrt(3/2) aint it since its further away from (0,2) to begin with

OpenStudy (anonymous):

ok i see my mistake \[D'(x)=4x^3-6x\]

OpenStudy (anonymous):

\[4x^3-6x=0\] \[2x(x^2-3)=0\] \[x=0, x=\pm\sqrt{\frac{3}{2}}\]

OpenStudy (anonymous):

i factored incorrectly ! lol

OpenStudy (amistre64):

:) factoring is for losers anyways, i say we always just go with the gut

OpenStudy (anonymous):

@amistre i still think distance formula is easier, but that is only because it is what i am used to. use the distance formula, take the derivative, set = 0 and solve. but i do like that you can do this without calc at all

OpenStudy (anonymous):

of course it helps if you solve correctly!

OpenStudy (amistre64):

well, i used calc to find the derivative to compare slope with; but that was all

OpenStudy (anonymous):

oh right. so no savings as far as calc goes.

OpenStudy (amistre64):

x^2 --> 2x , ow! i think i broke a sweat on that one ;)

OpenStudy (anonymous):

@wengzie in my own defense i did make several "it is early" disclaimers. method is right algebra is wrong @amistre, i differentiated a polynomial as well. it is the hidden machinery that is heavy, not the technique

OpenStudy (amistre64):

your method looks svelter tho, mine tends to drag knuckles. I do my math like a do my (......), theres alot of yelling and screaming, and in the end noone is happy

OpenStudy (anonymous):

spoken like a true mathematician!

OpenStudy (anonymous):

im quite confused now..hahaha...i'll follow ur flow of conversation ..:)))

OpenStudy (anonymous):

@AMISTRE64 and @SATELLITE73: You both are awesome!!! Thank you!!! I got it alright. :)